OpenStudy (anonymous):

y = 4x^2+2x-24 solve for finding the x intercept and vertex

hero (hero):

What the...???

OpenStudy (anonymous):

idk what im doing for it and i didn't think you were gonna help

OpenStudy (anonymous):

4x^2+2x-24 = 0 Solve for x for the intercepts

hero (hero):

Some of this you have to be able to do on your own.

hero (hero):

When I first started doing these nobody helped me

OpenStudy (anonymous):

same i just learned this last week

hero (hero):

They told me that all I had to do was FOIL and sent me out the door with pages of related hw problems

OpenStudy (anonymous):

im *

OpenStudy (anonymous):

x = 1/4 (-1-sqrt(97)) x = 1/4 (-1+sqrt(97))

OpenStudy (anonymous):

whats that? the vertex or ?

OpenStudy (anonymous):

add these two intercepts and then divided by 2 to find the x-coordinate of vertex